OpenStudy (anonymous):

Explain two ways in which you could simplify x5/x0. You may indicate an exponent in your answer with ^. For example, write x2 as x^2

OpenStudy (anonymous):

I think this is asking for the processes. One, you could apply the rule where you subtract exponents in a division problem. So 5-0=5, therefore f(x)=x^5/x^0=x^5. Also, because anything to the 0 power is 1 and since a number "n"/1=n, you know that f(x)=x^5

OpenStudy (anonymous):

thankk you (: